Make appropriate plots or perform linear regressio
Answer #1

Question a

part 1: zero order rate law

The integrated zero order rate law is as follows.

[A_t]=-kt+[A_0]

The graph between [At] and t should be linear.

1.4 1.2 0.8 0.6 0.4 0.2 8 10 12 0 4 Time

Part 2: First order rate law

The integrated first order rate law is as follows.

ln[A_t]=-kt+ln[A_0]

Therefore, a graph between ln [At] vs t must be linear if the reaction is first order.

0.4 0.2 12 0.2 0.4 -0.6 0.8 -1 1.2 1A

Part 3: Second Order rate law

The integrated second order rate law is as follows.

\frac{1}{[A_t]}=kt+\frac{1}{[A_0]}

Thus, the graph between 1/[At] vs t should be linear for a second order reaction.

3.5 0.242k+0 8032 2 = 0.9999 1.5 0.5 0 4 6 8 10 12

Thus, the highest linear regression was obtained for the second order graph (r2=0.9999). Thus, this is a second order equation.

QUESTION B

From the integrated rate law for the second order reaction, the rate constant k is given by the gradient of the graph.

\frac{1}{[A_t]}=kt+\frac{1}{[A_0]}

From the graph, the rate constant is,

k=0.242/(Ms)

QUESTION C

The half-life for the second order reactions is as follows.

t=\frac{1}{k[A_0]}

where [A0] is the initial concentration. Substituting,

t=\frac{1}{0.242/(Ms)*1.24M]}

t=0.300s

QUESTION D

Using the second order equation,

\frac{1}{[A_t]}=kt+\frac{1}{[A_0]}

Substituting,

\frac{1}{0.380M}=0.242/Ms*t+\frac{1}{1.24M}

t=7.54s